# Incentivizing supplemental math assignments and using AI-generated hints improve exam performance, especially for racially minoritized students

### v1

@@ IV Discussion
We offered two kinds of optional supports to students in an Introductory Physics course: supplemental math assignments and AI hints. We found that students who completed the math assignments and/or used AI hints performed better on the exams that were most related to these supports (Figs. 5 and 6).
Previous studies have shown that online supplemental math assignments can be effective in improving student performance in Physics courses [18, 4]. In this study, we tried a unique scaled incentive structure for free online supplemental math assignments to increase the completion rate, especially among students that need more support with mathematical skills. With this incentive structure, we were able to achieve higher completion rates among students who might need more support in the course compared to students who are well equipped to perform well in the course, as indicated by students’ incoming FCI scores. In other words, when incentivitized with the scaled extra credit, about 54 percent of students that scored below the class mean on the FCI completed the first supplemental math assignment compared to only 37 percent of students that scored at or above the class mean. By contrast, without the incentive, only 19 percent of students that scored below the class mean on FCI completed the first supplemental math assignment, compared to 24 percent of students that scored at or above the mean. When we broke this down by race/ethnicity, we found that our scaled incentive structure increased completion rate among students from historically under-represented racial/ethnic backgrounds. The higher supplemental math completion rate among PEERs, combined with the improved performance among students who completed supplemental math assignment 1, suggests that offering these assignments along with a scaled incentive structure can decrease racial inequities in Physics courses.
@@
Our study is one of the first to explore the impact of offering LLM (Large Language Model)-based AI generated hints during homework assignments on student performance in exams in an Undergraduate Physics course. Given the differences in quality of AI available in open-source compared to paid formats (GPT 3 vs 4 at the time of this study) [49, 50], differential access to AI could increase inequities in student performance on homework assignments and exams. In this course, high-quality AI hints for homework problems were made available to all students through the course platform Kudu. We found that a large majority of students used at least one AI hint and there were not significant differences in the use of AI hints by gender and race/ethnicity. Moreover, there was a positive association between use of AI hints and exam 2 performance, which was the exam most aligned with homework assignments. This suggests that high-quality generative AI tools, if made available to everyone, have the potential to support student learning equitably. In future studies, we plan to explore further students’ reasons for using or not using AI hints and their experience with using these hints.
One major limitation of our work is that despite the scaled extra credit incentive structure, the completion rate for supplemental math assignments is still less than 60 percent. Integrating the supplemental math assignments into formal class time as a component of discussion section might further increase student participation. Sharing the results of this study with future students may also be helpful in increasing the perceived utility value of the supplemental math assignments. Another limitation is that we were not able to disaggregate the race/ethnicity and gender data because of smaller sample sizes in some of the groups we have included in those categories. We recognize that combining these groups with different historical, social, and cultural context obscures patterns of racial and gender inequality [51]. In future work, we plan to collect larger datasets that could allow us to diasggregate these data for analyses.
In future studies, we also plan to explicitly measure student expectancies for performance in the course at the beginning and in the middle of the term. We also plan to collect some qualitative data using surveys and interviews to give us insights into the factors that shape students’ decisions to use the optional supports offered in the course. This qualitative data will help us understand the perceived value of these optional supports and inform structural modifications to increase student uptake of these supports.
@@ Acknowledgments
We thank Warren Essey, Stuart Brown, and Joshua Samani for valuable discussions and input. We also thank Harrison Parker for designing the teaching intervention flow chart. This work was supported by UCLA’s Teaching and Learning Center.
@@ Appendix B Sample AI-generated hints
Appendix C Other results in the study
@@
Homework score 1.43 0.27 $<0.001$
AI hints usage -0.18 0.20 0.356
Math supplement material 1 0.17 0.09 0.061
Math supplement material 2 -0.10 0.09 0.268
Force concept inventory 0.41 0.05 $<0.001$
Calculus concept inventory 0.20 0.05 $<0.001$
@@
Homework score 3.34 0.19 $<0.001$
AI hints usage 0.39 0.17 0.019
Math supplement material 1 0.08 0.08 0.320
Math supplement material 2 -0.02 0.08 0.824
Math supplement material 3 0.07 0.07 0.295
Force concept inventory 0.29 0.04 $<0.001$
Calculus concept inventory 0.15 0.04 $<0.001$
@@
Homework score 3.32 0.19 $<0.001$
Math supplement material 1 0.08 0.08 0.333
Math supplement material 2 -0.01 0.08 0.901
Math supplement material 3 0.07 0.07 0.331
Calculus concept inventory 0.15 0.04 $<0.001$
AI hints usage 0.33 0.17 0.051
Force concept inventory 0.33 0.04 $<0.001$
AI hints usage : Force concept -0.31 0.17 0.065

### v2

@@ IV Discussion
We offered two kinds of optional supports to students in an Introductory Physics course: supplemental math assignments and AI-generated hints. We found that completing the math assignments and/or using AI-generated hints was associated with better exam performance on the exams that were most related to these supports (Figs. 5 and 6).
Previous studies have shown that online supplemental math assignments can be effective in improving student performance in Physics courses [18, 4]. In this study, we tried a unique scaled incentive structure for free online supplemental math assignments to increase the completion rate, especially among students that need more support with mathematical skills. With this incentive structure, we were able to achieve higher completion rates among students who might need more support in the course compared to students who are well equipped to perform well in the course, as indicated by students’ incoming FCI scores. In other words, when incentivitized with the scaled extra credit, about 54 percent of students that scored below the class mean on the FCI completed the first supplemental math assignment compared to only 37 percent of students that scored at or above the class mean. By contrast, without the incentive, only 19 percent of students that scored below the class mean on FCI completed the first supplemental math assignment, compared to 24 percent of students that scored at or above the mean. When we broke this down by race/ethnicity, we found that our scaled incentive structure was associated with higher completion rate among students from historically under-represented racial/ethnic backgrounds. The higher supplemental math completion rate among PEERs, combined with the improved performance among students who completed supplemental math assignment 1, suggests that offering these assignments along with a scaled incentive structure can decrease racial inequities in Physics courses.
@@
Our study is one of the first to explore the impact of offering LLM (Large Language Model)-based AI-generated hints during homework assignments on student performance in exams in an Undergraduate Physics course. Given the differences in the quality of AI available in open-source compared to paid formats (GPT 3 vs 4 at the time of this study) [50, 51], differential access to AI could increase inequities in student performance on homework assignments and exams. In this course, high-quality AI-generated hints for homework problems were made available to all students through the course platform Kudu. We found that a large majority of students used at least one AI-generated hint and there were not significant differences in the use of AI-generated hints by gender and race/ethnicity. Moreover, there was a positive association between use of AI-generated hints and exam 2 performance, which was the exam most aligned with homework assignments. This suggests that high-quality generative AI tools, if made available to everyone, have the potential to support student learning equitably. In future studies, we plan to explore further students’ reasons for using or not using AI hints and their experience with using these hints.
IV.1 Limitations and Future Directions
Although our interventions are based on the Situated Expectancy-Value Theory of Achievement Motivation, we did not collect survey data on student’s expectancy beliefs and values regarding the supplemental math assignments and AI-generated hints. We also did not collect data on student use of any other AI-powered tools or use of any other optional supports such as tutoring services. We plan to address these shortcomings in future studies. Another limitation of our study is that we were not able to disaggregate the race/ethnicity and gender data because of smaller sample sizes in some of the groups we have included in those categories. We recognize that combining these groups with different historical, social, and cultural context obscures patterns of racial and gender inequality [52]. In future work, we plan to collect larger datasets that could allow us to diasggregate these data for analyses.
A limitation of our intervention is that despite the scaled extra credit incentive structure, the completion rate for supplemental math assignments is still less than 60 percent. Integrating the supplemental math assignments into formal class time as a component of discussion section might further increase student participation. Sharing the results of this study with future students may also be helpful in increasing the perceived utility value of the supplemental math assignments.
In future studies, we also plan to collect some qualitative data using surveys and interviews to give us insights into the factors that shape students’ decisions to use the optional supports offered in the course. This qualitative data will help us understand the perceived value of these optional supports and inform structural modifications to increase student uptake of these supports.
@@ Acknowledgments
We thank Warren Essey, Stuart Brown, and Joshua Samani for valuable discussions and input. We also thank Harrison Parker for designing the teaching intervention flow chart. This work was supported by UCLA’s Teaching and Learning Center and by the UCLA Department of Physics and Astronomy.
@@ Appendix B Sample AI-generated hints
Appendix C Sample isomorphic questions
In this section, we demonstrate a pair of exam and homework questions that are determined to be isomorphic in our exam question evaluation phase (Sec. II.3).
Homework problem: ”Block 2 slides along a horizontal frictionless table as block 1 falls. The blocks are connected via a frictionless pulley. Find the speed of the blocks after they have each moved 2.0 m. Assume that they start at rest and that the pulley has negligible mass. Use $m_{1}=2$ kg and $m_{2}=4$ kg.”
Exam problem: ”Block 2 with mass $m_{2}=3$ kg slides along a horizontal table as block 1 with mass $m_{1}=2$ kg moves downward. The blocks are connected by a rope via a frictionless pulley; the rope and the pulley have negligible mass. Find the speed of the blocks after they have each moved 2.0 m starting from rest.”
Appendix D Other results in the study
@@
Homework score 1.38 0.38 $<0.001$
AI hints usage -0.18 0.20 0.337
Math supplement material 1 0.13 0.08 0.122
Force concept inventory 0.42 0.05 $<0.001$
Calculus concept inventory 0.20 0.06 $<0.001$
@@
Homework score 3.34 0.32 $<0.001$
AI hints usage 0.40 0.14 0.004
Math supplement material 1 0.07 0.08 0.360
Math supplement material 3 0.07 0.08 0.372
Force concept inventory 0.29 0.05 $<0.001$
Calculus concept inventory 0.15 0.05 0.003
@@
Homework score 3.32 0.32 $<0.001$
Math supplement material 1 0.07 0.08 0.353
Math supplement material 3 0.07 0.08 0.399
Calculus concept inventory 0.15 0.05 0.004
AI hints usage 0.33 0.13 0.008
Force concept inventory 0.33 0.06 $<0.001$
AI hints usage : Force concept -0.31 0.18 0.094